# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

### v3

@@ Abstract
Despite the success of Instruction Tuning (IT) in training large language models (LLMs) to perform arbitrary user-specified tasks, these models often still leverage spurious or biased features learned from their training data, leading to undesired behaviours when deploying them in new contexts. In this work, we introduce Focus Instruction Tuning (FIT), which trains LLMs to condition their responses by focusing on specific features whilst ignoring others, leading to different behaviours based on what features are specified. Across several experimental settings, we show that focus-tuned models can be adaptively steered by focusing on different features at inference-time: for instance, robustness can be improved by focusing on core task features and ignoring spurious features, and social bias can be mitigated by ignoring demographic categories. Furthermore, FIT can steer behaviour in new contexts, generalising under distribution shift and to new unseen features at inference time, and thereby facilitating more robust, fair, and controllable LLM applications in real-world environments.
Machine Learning, ICML
@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next-token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023).
However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2023). This motivates the need for methods that can adaptively steer models at inference time, enabling users to control model behaviours based on evolving requirements or safety considerations.
@@
Accordingly, we propose Focus Instruction Tuning (FIT), an extension of traditional IT that fine-tunes LLMs with respect to an instruction specifying which features to “focus on” or “ignore.” FIT trains LLMs to condition responses based on these focus specifications and respond differently to the same task input based on the specified features, allowing end users to dynamically steer model behaviour. This capability provides precise, explainable control over features leveraged by models, and can be used to enforce desired invariances. For instance, in Figure 1, we illustrate how FIT can be used to steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
In our experiments, we demonstrate that FIT enables precise control over model behaviour by steering it to focus on task-relevant features while ignoring features specified as irrelevant or spurious (see Section 4). This flexibility allows FIT to address challenges such as mitigating the influence of demographic features in question-answering tasks, while also proving effective across diverse NLP tasks like sentiment analysis and natural language inference. Furthermore, we show that FIT is robust to distribution shifts over feature values and can generalise to new, unseen features, demonstrating its adaptability to dynamic user requirements and varying task contexts.
In summary, our primary contributions are as follows:
  1. 1.
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about core, spurious, or bias-relevant features in order to steer models according to the desired feature specification.
  2. 2.
    We experiment with FIT across several key NLP tasks, including sentiment analysis, natural language inference, and question-answering. We find that FIT is highly effective for steering behaviour on all tasks with respect to a variety of lexical, distributional, semantic, and demographic features.
  3. 3.
    We show that focus tuning generalizes both to new features not seen during training, and to distribution shift over feature values.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold (Wang et al., 2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Arjovsky et al., 2019; Xiao et al., 2021; Venkataramani et al., 2024; Hemmat et al., 2024), texture (Geirhos et al., 2018; Baker et al., 2018), or scene elements (Hemmat et al., 2024), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift. Traditional approaches for detecting and mitigating spurious feature learning, particularly under distribution shifts, include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), counterfactual inference (Wang & Culotta, 2020, 2021; Udomcharoenchaikit et al., 2022), or generating synthetic interventional data (Bansal & Grover, 2023; Yuan et al., 2024; Wang et al., 2024).
@@
Substantial work in mechanistic interpretability has also aimed to discover models’ latent representation of, and reliance on, various features (Davies & Khakzar, 2024). For instance, causal probing trains supervised probing classifiers to predict and modify feature representations encoded by foundation models Belinkov, 2022, and has been deployed to study how LLMs leverage task-causal versus spurious features (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023; Canby et al., 2024). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020, 2022, 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Instruction Tuning. Due to the next-word prediction training objective, large language models (LLMs) often struggle by default to generate outputs that align with human instructions in downstream applications (Huang et al., 2023). Instruction-tuning (IT) mitigates this issue by fine-tuning pre-trained LLMs on datasets composed of instruction-response pairs (Zhang et al., 2023), aiming to align the responses of the fine-tuned model more closely with the distributions preferred by humans (Ouyang et al., 2022). There are several popular approaches for collecting IT training data, such as using human-annotated data (Dolly, 2023), extracting datasets from existing collections (Longpre et al., 2023; Mishra et al., 2022), or gathering data from internet sources (Zhou et al., 2024). IT datasets can also be synthesised with LLMs, either by bootstraping them from the same model that will be instruction-tuned on them (Wang et al., 2023c; Chen et al., 2024), or by distilling from a larger or more powerful model to instruction-tune smaller models (Taori et al., 2023; Mitra et al., 2023; Xu et al., 2023).
Despite the success of IT in zero-shot generalisation, Gudibande et al. (2023) find that improvements on many downstream benchmark tasks may be largely due to coverage of task data within IT training datasets; and bootstrapping IT methods (which, in principle, might not be subject to this issue provided they synthesise novel IT task instances) require a robust and effective LLM for fine-tuning to avoid degenerate training cycles (Zhang et al., 2023). Furthermore, Kung & Peng (2023) show that some of the downstream performance gains from IT can be attributed to models’ ability to learn surface-level patterns, such as the required answer format, rather than acquiring more generalisable instruction-following skills. These limitations underscore the need for advancements in supervised fine-tuning (SFT) methods beyond IT to facilitate more predictable and reliable control of downstream model behaviours.
@@
Aligning LLMs. Alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) are powerful tools for aligning LLMs with annotated preference data and lead to reduced prevalence of harmful behaviour (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023; Korbak et al., 2023). However, RLHF-trained models still exhibit key alignment limitations such as sycophancy (defaulting to agreement with users even when incorrect or harmful; Perez et al., 2023; Sharma et al., 2024), and can still be adversarially prompted to generate harmful responses (Carlini et al., 2024). Furthermore, even well-aligned models can rapidly fall out of alignment when they are fine-tuned (Zhan et al., 2024; Yang et al., 2024; Lermen & Rogers-Smith, 2024), even on benign tasks (Qi et al., 2023). Thus, effectively steering model behaviours across the range of potential safety concerns that might emerge during LLM pre-training, fine-tuning, or deployment to novel contexts is an important and challenging goal in AI safety, necessitating more flexible and generalisable steering methods (Anwar et al., 2024).
@@
A growing literature has worked to address this challenge via inference-time steering, where LLMs do not need to be retrained with respect to safety limitations, but can instead be controlled at inference time to steer models towards desirable behaviours or away from undesirable ones. For instance, latent steering methods perform embedding-space interventions that push models towards desirable behaviours or away from undesirable ones (Turner et al., 2023; Zou et al., 2023; Bhattacharjee et al., 2024; Li et al., 2024; Han et al., 2024). However, these methods require white-box access to model representations at inference time; and interventions are specific to each target behaviour, meaning that they must be re-computed in the context of any given target behaviour. Our work addresses these limitations by training models to control and adapt their responses based on explicit focus specifications, enabling users to flexibly and dynamically steer model behaviour at inference-time using simple, natural-language instructions.
@@ 3 Methodology
Preliminaries. We consider a pre-trained, decoder-only large language model (LLM) $p_{\theta}$ that models the probability of token sequences autoregressively over its vocabulary $\mathcal{V}$. Given a sequence of tokens $s=[s_{1},\dots,s_{L}]\in\mathcal{V}^{L}$, the jonit probability of $s$ under the model is given as
$p_{\theta}(s)\;=\;\prod_{i=1}^{L}\,p_{\theta}(s_{i}\mid s_{<i}),$ (1)
where $p_{\theta}(s_{1}\mid\emptyset)=p_{\theta}(s_{1})$. In supervised fine-tuning, we minimise the negative log-likelihood (NLL) of the output tokens $y$ given the input tokens $x$ using the autoregressive formulation defined in Equation 1.
In instruction tuning (IT) (Zhang et al., 2023), a form of SFT, an additional task instruction $I$ accompanies the input-output pair $(x,y)$, forming a tuple $(I,x,y)$. The objective becomes minimising the NLL of $y$ given both $I$ and $x$ over the distribution of input-output pairs, and instructions.
Focus Instruction Tuning (FIT). We introduce Focus Instruction Tuning (FIT), a specialised form of instruction tuning that trains LLMs to adjust their responses based on user-specified features provided in natural language.
Let $\mathcal{F}$ denote the set of possible features (e.g., specific keywords, sentiment, verb tense, demographic information, etc.) that the model can be instructed to focus on or ignore when generating responses.
We consider a set of natural language instructions to focus or rule out specified features in $\mathcal{F}$ which we term the focus instruction set $\mathcal{I}_{\text{focus}}$. Explicitly, we define $\mathcal{I}_{\text{focus}}$ as
@@
$,\;{\color[rgb]{0,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62109375,0.44921875}\text{focus}(F_{i})\wedge\text{ignore}(F_{j})}\mid F_{i},F_{j}\in\mathcal{F}\},$
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT. For specific examples of focus instructions that we consider, see Appendix E.
@@
Consider a sample $(x,y)\sim p_{\text{data}}$ drawn from an underlying data distribution, and a focus instruction $I_{\text{focus}}$ drawn from a distribution $p_{\mathcal{I}_{\text{focus}}}$ over the set of focus instructions $\mathcal{I}_{\text{focus}}$. Then the likelihood of response $y$ conditioned on input $x$, task instruction $I$ (as in standard IT), and focus-instruction $I_{\text{focus}}$ is modelled as $p_{\theta}(y|I,I_{\text{focus}},x)$.
@@
FIT Training. Consider a classification task with finite label space $\mathcal{Y}$, where a single core feature $C\in\mathcal{F}$ is fully predictive of label $y\in\mathcal{Y}$ given input $x$ at both training time and under distribution shift (Koh et al., 2021). We also consider spurious features $S\in\mathcal{S}\subseteq\mathcal{F}$ from a subset of spurious features $\mathcal{S}$, where feature values $s\in\text{Val}(S)$ for some spurious feature $S\in\mathcal{S}$ correlate with a label $y_{s}\in\mathcal{Y}$, where this correlation may change under distribution shift (Ming et al., 2022). Finally, we define $\mathcal{F}$ as the set of features that may be included in focus instructions during training, consisting of the core feature and the set of spurious features =$\mathcal{F}=\{C\}\cup\mathcal{S}$.
For a sample $(x,y)\sim p_{\text{data}}$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)\in\mathcal{Y}$ that depends on the ground truth label $y$, focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ and spurious feature value $s\in\text{Val}(S)$ present in $x$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying core feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the core feature is ignored, $y_{\text{focus}}$ is defined as $y_{\text{focus}}=y_{s}$, where $y_{s}$ is the label spuriously correlated with the particular spurious feature value $s$ present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on feature specifications implemented through focus instructions. More formally, we define $y_{\text{focus}}$ as:
$y_{\text{focus}}=\begin{cases}y&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{C};\\ y_{s}&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{s}\;,\end{cases}$ (3)
where we define the task-casual and spurious instruction target sets as
$\mathcal{I}_{\text{focus}}^{y}$ =$=\{\emptyset,\;\text{focus}(C),\;\text{focus}(C)\wedge\text{ignore}(S),\;$ (4)
$\qquad\qquad\qquad\qquad\;\;\text{ignore}(S)\mid S\in\mathcal{S}\},$ (5)
$\mathcal{I}_{\text{focus}}^{y_{s}}$ =$=\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j}),$ (6)
$\ \qquad\qquad\qquad\qquad\mid F_{j}\in\mathcal{F}\setminus\{S\}\},$ (7)
respectively. See Figure 2 for a concrete example showing the focus label values for an example from the MNLI dataset under different focus instructions.
@@
The objective of FIT training is to minimise the negative log-likelihood (NLL) of the response $y_{\text{focus}}$ conditioned on $I,I_{\text{focus}},x$. Formally, as a form of expected-risk minimisation (ERM) (Vapnik et al., 1998), we define the FT loss objective as:
$\min_{\theta}\mathbb{E}_{(x,y)\sim p_{\text{data}},\;I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}}\;\left[-\log p_{\theta}\left(y_{\text{focus}}\mid I,I_{\text{focus}},x\right)\right].$ (8)

### v4

@@ Abstract
Despite the success of Instruction Tuning (IT) in training large language models (LLMs), such models often leverage spurious or biased features learnt from their training data and can become misaligned, leading to undesired behaviours. While existing techniques can steer model behaviour at inference-time, they are often post-hoc and do not embed steering as an intrinsic model feature. In this work, we introduce Focus Instruction Tuning (FIT), which trains LLMs to condition their responses by focusing on specific features whilst ignoring others, leading to different behaviours based on what features are specified. Across diverse benchmarks, we demonstrate that FIT: (i) successfully steers behaviour at inference time; (ii) increases robustness by amplifying core task signals and down-weighting spurious cues; (iii) mitigates social bias by suppressing demographic attributes; and (iv) generalises under distribution shifts and to previously unseen focus features. FIT therefore offers a lightweight, intrinsic mechanism for building more robust, fair, and easily controllable LLMs.111 Our project page, including links to codebase and datasets, is available at: https://tomalamb.github.io/focus-instruction-tuning/.
Machine Learning, ICML, Instruction Tuning, LLMs, Spurious Correlations
@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised supervised fine-tuning approach, has become integral to developing large language models (LLMs) capable of effectively following instructions (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training on next-token prediction allows LLMs to learn broad patterns and knowledge, IT further fine-tunes models using input-output pairs accompanied by natural-language instructions, enhancing their ability to handle diverse, open-ended tasks (Huang et al., 2023).
@@
Despite notable gains in zero-shot generalisation from IT, recent studies indicate that these improvements may be superficial, primarily due to models simply learning task formats or spurious correlations rather than developing genuine understanding or more generalisable instruction-following capabilities (Kung & Peng, 2023; Ghosh et al., 2024). Consequently, models often fail in new contexts lacking these patterns (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2024). Existing representation-level interventions aiming to steer model behaviour at inference-time to overcome issues such as misalignment serve as post hoc corrections, becoming increasingly complex and impractical with modern large-scale models (Bhattacharjee et al., 2024; Li et al., 2024). This underscores the necessity for simple, intrinsic methods enabling dynamic steerability to align model behaviour with evolving user and safety needs.
To address this, we propose Focus Instruction Tuning (FIT), an extension of IT that fine-tunes LLMs with respect to instructions specifying which features to focus on” or ignore.” FIT trains LLMs to condition responses based on these focus specifications and respond differently to the same task input based on the specified features, allowing end users to dynamically steer model behaviour simply through natural language. This capability provides precise, explainable control over features leveraged by models, and can be used to enforce desired invariances. For instance, in Figure 1, we illustrate how FIT can be used to steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
Our experiments demonstrate that FIT precisely steers models to emphasise task-relevant features while disregarding irrelevant or spurious ones, effectively mitigating biases. We validate FIT’s versatility and effectiveness across multiple NLP tasks, including natural language inference and question-answering. Additionally, FIT robustly generalises under distribution shifts and to unseen features, underscoring its adaptability.
In summary, our primary contributions are:222 Individual contributions of each author are listed in Appendix A.
  1. (a)
    Dynamic Steerability. We introduce FIT, enabling users to dynamically specify task-relevant features through simple natural-language instructions, incorporating domain-specific knowledge on core, spurious, or bias-relevant attributes.
  2. (b)
    Broad Task Effectiveness. We validate FIT’s effectiveness across diverse NLP tasks, including sentiment analysis, natural language inference, and question-answering, demonstrating precise control over lexical, distributional, semantic, and demographic features.
  3. (c)
    Robust Generalisation. We show that FIT generalises robustly both to unseen features during training and under distributional shifts in feature values.
  4. (d)
    Preservation of Core Capabilities. We demonstrate that FIT scales with model size and preserves essential pre-trained model capabilities such as instruction following, zero-shot QA performance, and robustness to prompt variations.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset i.e., input features that are correlated with outputs in the training distribution, but which are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold (Wang et al., 2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Arjovsky et al., 2019; Xiao et al., 2021; Venkataramani et al., 2024; Hemmat et al., 2024), texture (Geirhos et al., 2018; Baker et al., 2018), or scene elements (Hemmat et al., 2024), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token “SPIELBERG” highly co-occurs with positive sentiment in SST-2 (Socher et al., 2013b; Wang & Culotta, 2020), a binary sentiment analysis dataset, meaning that models trained on SST-2 may learn to predict sentiment by leveraging this feature as a spurious feature instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-task-causal features undermines the robustness of models when generalising under distribution shift.
Traditional approaches for detecting and mitigating spurious feature learning, particularly under distribution shifts, include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), counterfactual inference (Wang & Culotta, 2020, 2021; Udomcharoenchaikit et al., 2022), or generating synthetic interventional data (Bansal & Grover, 2023; Yuan et al., 2024; Wang et al., 2024). Other recent work, such as conditional supervised fine-tuning (cSFT), aims to mitigate spurious correlations by conditioning training on feature labels, effectively discouraging the model from relying on dataset-specific biases (Zhang et al., 2024). However, cSFT does not support dynamic adaptation to new spurious features at test time, which may emerge due to distribution shifts or arise from misalignment introduced during further stages of training (Zhan et al., 2024; Zhou et al., 2024b).
@@
Substantial work in mechanistic interpretability has also aimed to discover models’ latent representation of, and reliance on, various features. For instance, causal probing involves training supervised probing classifiers to predict and modify feature representations encoded by foundation models (Belinkov, 2022; Davies & Khakzar, 2024), and has been deployed to study how LLMs leverage task-causal versus spurious features (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023; Canby et al., 2024). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating the use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, in order to remove the use of biased features concept removal methods aim to locate and manipulate the corresponding supervised representations encoded by foundation models (Ravfogel et al., 2020, 2022, 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Instruction Tuning. Due to the next-word prediction training objective, foundation language models often struggle by default to generate outputs that align with human instructions in downstream applications (Huang et al., 2023). Instruction-tuning (IT) mitigates this issue by fine-tuning pre-trained LLMs on datasets composed of instruction-response pairs (Zhang et al., 2023), aiming to align the responses of the fine-tuned model more closely with the distributions preferred by humans (Ouyang et al., 2022). There are several popular approaches for collecting IT training data, such as using human-annotated data (Dolly, 2023), extracting datasets from existing collections (Longpre et al., 2023; Mishra et al., 2022), or gathering data from internet sources (Zhou et al., 2024a). IT datasets can also be synthesised with LLMs, either by bootstrapping them from the same model that will be instruction-tuned (Wang et al., 2023c; Chen et al., 2024), or by distilling from a larger or more powerful model to instruction-tune smaller models (Taori et al., 2023; Mitra et al., 2023; Xu et al., 2023).
Despite the success of IT in zero-shot generalisation, recent findings indicate that downstream performance improvements from IT often arise due to models learning surface-level patterns, such as specific answer formats, rather than genuinely acquiring generalisable instruction-following skills (Kung & Peng, 2023; Ghosh et al., 2024; Zhou et al., 2023). Moreover, it has been demonstrated IT performance gains frequently come at a cost, referred to as the “alignment tax” (Ouyang et al., 2022), whereby models exhibit enhanced instruction-following capabilities but suffer performance degradation on other standard task benchmarks (Ouyang et al., 2022; Ren et al., 2024). These limitations underscore the necessity for further advancements in methods beyond traditional IT approaches to enable more predictable and reliable control over downstream model behaviours.
@@
Aligning LLMs. Alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) are powerful tools for aligning LLMs with annotated preference data and lead to reduced prevalence of harmful behaviour (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023; Korbak et al., 2023). However, RLHF-trained models still exhibit key alignment limitations such as sycophancy (defaulting to agreement with users even when incorrect or harmful; Perez et al., 2023; Sharma et al., 2024), and can still be adversarially prompted to generate harmful responses (Carlini et al., 2024). Furthermore, even well-aligned models can rapidly fall out of alignment when they are fine-tuned (Zhan et al., 2024; Yang et al., 2024b; Lermen & Rogers-Smith, 2024), even on benign tasks (Qi et al., 2024). Methods such as SteerLM (Dong et al., 2023) use fixed, human-annotated stylistic attributes and iterative bootstrapping, focusing on output style; however, their reliance on predefined attributes and control tokens limits flexibility, adaptability, and generalisation to unseen features. Thus, effectively steering behaviours across the diverse range of environments foundation models will experience during their training and deployment remains an important and challenging goal in AI safety, necessitating more flexible and generalisable steering methods (Anwar et al., 2024).
@@
A growing literature has worked to address the challenge of misalignment via inference-time steering, where LLMs do not need to be retrained with respect to safety limitations, but can instead be controlled at inference time to steer models towards desirable behaviours or away from undesirable ones. For instance, latent steering methods (LSMs) perform embedding-space interventions that push models towards desirable behaviours or away from undesirable ones (Turner et al., 2023; Zou et al., 2023; Bhattacharjee et al., 2024; Li et al., 2024; Han et al., 2024). However, these methods require white-box access to model representations during inference, and interventions must be computed separately for each target behaviour. Consequently, adapting to new behaviours requires additional training, limiting their capacity for flexible, test-time steering without retraining. Thus, existing methods remain fundamentally post-hoc solutions, failing to embed steerability intrinsically within models as a generalisable capability.
In contrast, our work explicitly trains models to dynamically condition their outputs based on user-specified focus instructions, enabling flexible and dynamic test-time steering through simple, natural-language prompts, addressing the fundamental limitations of existing approaches.
@@ 3 Methodology
3.1 Preliminaries
We consider a pre-trained, decoder-only LLM, $p_{\theta}$, that models the probability of token sequences autoregressively over its vocabulary $\mathcal{V}$. Given a sequence of tokens $s=(s_{1},\dots,s_{L})\in\mathcal{V}^{L}$, the joint probability of $s$ under the model is given as
$p_{\theta}(\bm{s})\;=\;\prod_{i=1}^{L}\,p_{\theta}(s_{i}\mid s_{<i}),\quad y_{<i}=(y_{1},\cdots,y_{i-1}),$ (1)
where $p_{\theta}(s_{1}\mid\emptyset)=p_{\theta}(s_{1})$. In supervised fine-tuning (SFT), we minimise the negative log-likelihood (NLL) of output sequences $y\in\mathcal{V}^{\mid y\mid}$ given input sequences $x\in\mathcal{V}^{\mid x\mid}$ using the autoregressive formulation defined in Equation 1.
In IT (Zhang et al., 2023), a form of SFT, an additional task instruction $I\in\mathcal{V}^{\mid I\mid}$ accompanies the input-output sequence pair forming a tuple $(I,x,y)\in\mathcal{V}^{\mid I\mid\times\mid x\mid\times\mid y\mid}$. The objective becomes the minimisation of the expected NLL of $y$ given both $I$ and $x$ over the distribution of input-output pairs and instructions.
3.2 Focus Instruction Tuning (FIT)
We introduce Focus Instruction Tuning (FIT), a specialised form of instruction tuning that trains LLMs to adjust their responses based on user-specified features provided in natural language.
Focus Instructions. Let $\mathcal{F}$ denote the set of possible features (e.g., specific keywords, sentiment, verb tense, demographic information, etc.) that the model can be instructed to focus on or ignore when generating responses. We consider a set of natural language instructions to focus or rule out specified features in $\mathcal{F}$ which we term the focus instruction set $\mathcal{I}_{\text{focus}}$. Explicitly, we define $\mathcal{I}_{\text{focus}}$ as
@@ 3.2 Focus Instruction Tuning (FIT)
$,\;{\color[rgb]{0,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.46875,0.33203125}\text{focus}(F_{i})\wedge\text{ignore}(F_{j})}\mid F_{i},F_{j}\in\mathcal{F}\},$
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt during training to help the model learn both the underlying task and how to dynamically refocus its attention on user-specified features during FIT. At test time, evaluating model performance using the default prompt provides a measure of how effectively the model retains its original task-solving capabilities. For specific examples of the focus instructions considered, see Appendix C.
@@
Focus Labels.
@@
Consider a classification task with a finite label space $\mathcal{Y}$. A single core feature $C\in\mathcal{F}$ is fully predictive of the label $y\in\mathcal{Y}$ for any input $x$ at both training time and under distribution shift (Koh et al., 2021). In addition, we have a subset of spurious features $\mathcal{S}\subseteq\mathcal{F}$. For each spurious feature $S\in\mathcal{S}$, values $s\in\operatorname{Val}(S)$ correlate with a label $y_{s}\in\mathcal{Y}$, where this correlation may change under distribution shift (Ming et al., 2022). Altogether, the set of features that can appear in focus instructions is =$\mathcal{F}=\{C\}\cup\mathcal{S}$.
For a sample $(x,y)\sim p_{\text{data}}$, we define the focus label
$y_{\text{focus}}\!\bigl{(}I_{\text{focus}},s,y\bigr{)}\;\in\;\mathcal{Y},$
which depends on the original ground-truth label $y$, the focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$, and the specific spurious feature value $s\in\operatorname{Val}(S)$ present in $x$. Intuitively, the focus label equals the ground-truth label ($y_{\text{focus}}=y$) when no focus features are specified (empty instruction $\emptyset$), when focusing on the core feature $C$, or when explicitly ignoring a spurious feature $S$. Conversely, when the instruction explicitly targets a spurious feature, we set $y_{\text{focus}}=y_{s}$, the label spuriously correlated with the concrete spurious value $s$ in $x$. Using the $y_{\text{focus}}$ labels as targets during training teaches the model to adapt its outputs to the feature specifications given in the focus instruction. See Figure 2 for a concrete illustration, and Definition 1 for a formal definition.
Definition 1 (Focus Labels).
For a sample $(x,y)\sim p_{\text{data}}$ and a focus instruction $I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}$, we define $y_{\text{focus}}=y_{\text{focus}}\!\bigl{(}I_{\text{focus}},s,y\bigr{)}$ for a spurious feature value $s\in\text{Val}(S)$ present in $x$ as:
$y_{\text{focus}}=\begin{cases}y&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{c},\\ y_{s}&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{s}\;,\end{cases}$
where the core and spurious instruction target sets are given as
$\mathcal{I}_{\text{focus}}^{c}$ =$=\{\emptyset,\;\text{focus}(C),\;\text{focus}(C)\wedge\text{ignore}(S)\mid\text{ignore}(S)\},$
$\mathcal{I}_{\text{focus}}^{s}$ =$=\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j})\mid F_{j}\in\mathcal{F}\setminus\{S\}\},$
respectively.
In summary, focus labels for instructions in $\mathcal{I}_{\text{focus}}^{c}$ coincide with the ground-truth label, since the focus is on the core feature, whereas focus labels for $\mathcal{I}_{\text{focus}}^{s}$ are the spurious labels associated with each spurious feature value. Refer again to Figure 2 for a worked example.
@@
FIT Training Objective. The objective of FIT training is to minimise the expected negative log-likelihood (NLL) of the response $y_{\text{focus}}$ conditioned on $I,I_{\text{focus}},x$. Formally, as a form of expected-risk minimisation (ERM) (Vapnik et al., 1998), writing $(x,y)\sim p_{\text{data}}$ and $I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}$, we define the FIT loss objective as:
$\min_{\theta}\mathbb{E}_{x,y,I,I_{\text{focus}}}\left[-\log p_{\theta}\left(y_{\text{focus}}\mid I,I_{\text{focus}},x\right)\right].$ (3)